OpenStudy (gabylovesyou):

Identify the correct slope and y intercept of the equation 2x + 4y = 12.

OpenStudy (anonymous):

Can you put the equation into slope intercept form? y = mx + b

OpenStudy (gabylovesyou):

slope = 2; y intercept at (0, 12) slope = 1/2; y intercept at (0, 3) slope = 1/2; y intercept at (0, 4) slope = -1/2); y intercept at (0, 3)

OpenStudy (gabylovesyou):

idk how to @qpHalcy0n

OpenStudy (anonymous):

Well, we need to get the y by itself. Start by subtracting 2x from both sides.

OpenStudy (gabylovesyou):

the answer is D!

OpenStudy (gabylovesyou):

right??

OpenStudy (anonymous):

You got it.